let f(x)={3x^2,x<=1}{ax+b,x>1} Find the value of a and b so that f is differentiable at x =1


1
Expert's answer
2020-09-06T18:43:30-0400

"f(x) = \begin{cases}
 3x^{2} &\text{if } x<=1 \\
 ax+b &\text{if } x>1
\end{cases}
\\\lim\limits_{x\mapsto1}ax+b=3\implies a+b=3
\\ \tfrac{d}{dx}(3x^{2})\mid_{x=1}=\tfrac{d}{dx}(ax+b)\mid_{x=1}\implies a=6\implies b=-3
\\Answer:a=6, b=-3"